# LOGIN: A Large Language Model Consulted Graph Neural Network Training Framework

### v1

@@ Abstract.
Large Language Model, Graph Neural Network, LLM, GNN
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Data miningccs: Information systems Information systems applications
@@ 1. Introduction
Individual entities along with their respective interactions, typically denoting as $\langle{u},e_{u,v},{v}\rangle$, where $u$ and $v$ are entities and $e_{u,v}$ is the interaction between $u$ and $v$, can be effectively represented and analyzed in the form of graph-structured data. For instance, reciprocal following in social networks (Golbeck and Hendler, 2006; Mccallum et al., 2007), financial transfers between accounts (Akoglu et al., 2014; Pourhabibi et al., 2020), chemical bonds between molecules (Blundell, 1996; Kearnes et al., 2016), and citations of academic articles (Leskovec et al., 2005; Alcacer and Gittelman, 2006) can be generally represented as graph-like structures. It thus burgeons graph machine learning for various tasks in data mining-related fields for decades, such as recommendation (Ying et al., 2018; Ma et al., 2019; Basu et al., 1998), anomaly detection (Liu et al., 2021a; Zhong et al., 2020; Krügel et al., 2002), drug discovery (Stärk et al., 2022; Wang et al., 2023b; Lo et al., 2018), and etc (Hamilton et al., 2017b; Riesen and Bunke, 2008; Broder et al., 2000). Recently, thanks to the advent of graph neural networks (GNNs) (Kipf and Welling, 2016; Hamilton et al., 2017a; Velickovic et al., 2017) and their demonstrated extensive capabilities, graph machine learning has become a highly active field attracting significant research attention. More than one-third of accepted papers in the research track of KDD’23 explicitly contain the term “graph” in their titles, underscoring the significant popularity of this topic [^1] https://kdd.org/kdd2023/research-track-papers/.
@@
Among these existing works, most of them strive to design different architectures of GNN to adapt to distinct graph types. The reason for that is the broad real-world sources of graph data induce considerable diversity across different graphs. One of the basic categorizations, for example, could be the case of homophily and heterophily (Rogers and Bhowmik, 1970a, b). Traditional GNNs, due to their fundamental mechanisms of message passing and aggregation (Battaglia et al., 2018), have demonstrated superior performance on homophilic graphs but fail to generalize to heterophily scenarios where dissimilar nodes are connected (Zhu et al., 2020; Luan et al., 2020; Liu et al., 2022). Consequently, researchers have committed to developing various structures, aiming at improving the effectiveness of GNNs on particular types of graphs (Lim et al., 2021; Liu et al., 2022; Pei et al., 2019; Luan et al., 2020). In addition to individualized designs, there are also observed recent works applying Graph Neural Architecture Search (GNAS) techniques for seeking better GNN architecture from a series of essential components (Zhou et al., 2022; Zhang et al., 2023). Despite their various routes of the existing works, these methods hold a similar methodology: Design a specialized style of graph neural networks to tackle specific analysis tasks.
@@
In this paper, our objective is to streamline the GNN design process and answer the question of whether diverse graph-structured data can be addressed using a simple and unified architecture. To this end, a powerful off-shelf tool, the Large Lanuange Model (LLM), is adopted to help us achieve the goal. LLMs have not only achieved remarkable advancements in natural language processing (Brown et al., 2020; Touvron et al., 2023), but have also buoyed promising insights in other domains, such as computer vision (Alayrac et al., 2022; Liu et al., 2023b) and recommendation systems (Ren et al., 2023; Mysore et al., 2023). This success can be attributed to the wealth of open-world knowledge stored in their large-scale parameters (Zhao et al., 2023a; Pan et al., 2024) and the emerging capabilities in logical reasoning (Huang and Chang, 2022; Brown et al., 2020).
Analogously, we aim to leverage the advantages of LLMs to enhance graph machine learning. Some pioneering research has explored the potential of LLMs on graph data, which can be typically categorized into LLMs-as-Predictors and LLMs-as-Enhancers, as shown in Fig. 1 (a) and (b). For the first kind, taking the node classification task as an example, some researchers harness the zero-shot and few-shot reasoning ability of LLMs to directly obtain node labels (Zhao et al., 2023b; Fatemi et al., 2023; Wang et al., 2023a; Guo et al., 2023). Others apply instruction tuning or prefix tuning to adapt LLMs specifically for graph tasks (Chai et al., 2023; Ye et al., 2023; Tang et al., 2023). However, most of them disregard the advantages of GNNs, and tuning LLMs may require substantial computing resources as well. For the latter one, LLMs-as-Enhancers, the LLMs are commonly utilized to enhance nodes’ semantic features (Chen et al., 2023; He et al., 2023) or to refine local topological structures (Sun et al., 2023). This kind of route is straightforward to apply LLMs in data preprocessing as a one-time enhancer, failing to couple the LLMs and GNNs interactively.
Different from the existing work, we formulate a new paradigm that integrates LLMs with GNNs in an interactive manner, and we coin it “LLMs-as-Consultants” (c.f. Fig. 1 (c)). The core difference resides in the insight that LLMs may explicitly contribute to the training process of GNN.
@@
Under this paradigm, we propose a framework named L O G IN , short for L LM c O nsulted G NN tra IN ing , empowering the interactive utilization of LLMs within the GNN training process. As an interactive approach, the crucial issues of L O G IN lie in what GNNs should deliver to LLMs and how to feed the LLMs’ responses back to GNNs. First, for LLMs’ inputs, we delve into prompt engineering to craft concise prompts for spotted nodes, which carry comprehensive semantic and topological information. Second, to utilize the responses from LLMs, we devise a complementary coping mechanism depending on their correctness. Specifically, compared to ground truth labels, when LLMs predict correctly, we update node features to obtain semantic enhancement. Otherwise, with the awareness of the sycophancy of LLMs, we impute the misclassification to the potential presence of local topological noises, hence performing structure refinement. Besides, particular criteria may serve in the selection of essential nodes to reduce time complexity. In our implementation, we adopt GNN predictive uncertainty to assess the necessity of consulting LLMs.
@@
To demonstrate the versatility and applicability of the proposed framework, we explore the effectiveness of L O G IN on node classification tasks across both homophilic and heterophilic graphs. By empirical studies, we illustrate that even basic baselines, when employed within the proposed LLMs-as-Consultants paradigm, can achieve comparable performance to advanced GNNs with intricate designs.
Figure 1. The paradigms for integrating LLMs with graphs.
@@ 2.1. Text-Attributed-Graphs (TAGs)
where $\mathcal{V}$ is a set of $N$ nodes, $\mathbf{A}\in\{0,1\}^{N\times N}$ denotes the adjacency matrix, and $s_{n}\in\mathcal{D}^{L_{n}}$ is the sequential text attached to node $n\in\mathcal{V}$, with $\mathcal{D}$ as the word dictionary, and $L_{n}$ as the sequence length. Note that our proposed framework is not limited to traditional TAGs that literally have texts as original attributes. In fact, most entities and their relations can be modeled and processed as graphs, and their characteristics can be expressed in text form.
@@ 2.2. Graph Neural Networks (GNNs)
For GNN training, the texts associated with the nodes should be encoded to an embedded space. We represent the node embeddings as $\mathbf{X}\in\mathbb{R}^{N\times D}$, in which each row $\boldsymbol{}{x}_{n}$ denotes the corresponding node embedding, with $D$ as its dimension number. For node classification, GNNs aggregate information from a node’s neighbors, and then update the node representation with the aggregated information. The k-th layer of a GNN can be formalized as:
(2) $x_{i}^{(l)}=f^{(l)}((\text{AGG}_{j\in\mathcal{N}(i)}^{(l)}x_{j}^{(l-1)}),x_{i}^{(l-1)}),$
where $x_{i}^{(l)}\in\mathbb{R}^{D}$ denotes the representation of the $i$-th node in the $l$-th layer representation, with $\mathcal{N}(i)$ as its neighborhood. In addition, $\text{AGG}^{(l)}$ is a function operator that is differentiable, and $f^{(l)}$ represents the certain structure of the $l$-th layer GNN.

### v2

@@ 1. Introduction
Individual entities along with their respective interactions, typically denoting as $\langle{u},e_{u,v},{v}\rangle$, where $u$ and $v$ are entities and $e_{u,v}$ is the interaction between $u$ and $v$, can be effectively represented and analyzed in the form of graph-structured data. For instance, reciprocal following in social networks (Golbeck and Hendler, 2006; Mccallum et al., 2007), financial transfers between accounts (Akoglu et al., 2014; Pourhabibi et al., 2020), chemical bonds between molecules (Blundell, 1996; Kearnes et al., 2016), and citations of academic articles (Leskovec et al., 2005; Alcacer and Gittelman, 2006) can be generally represented as graph-like structures. It thus burgeons graph machine learning in various tasks for decades, such as recommendation (Ying et al., 2018; Ma et al., 2019; Basu et al., 1998), anomaly detection (Liu et al., 2021; Zhong et al., 2020; Krügel et al., 2002), drug discovery (Stärk et al., 2022; Wang et al., 2023b; Lo et al., 2018), and etc (Hamilton et al., 2017b; Riesen and Bunke, 2008; Broder et al., 2000). Recently, thanks to the advent of graph neural networks (GNNs) (Kipf and Welling, 2016; Hamilton et al., 2017a; Velickovic et al., 2017) and their demonstrated extensive capabilities, graph machine learning has become a highly active field attracting significant research attention.
@@
Among these existing works, most of them strive to design different architectures of GNN to adapt to distinct graph types. The reason for that is the broad real-world sources of graph data induce considerable diversity across different graphs. One of the basic categorizations, for example, could be the case of homophily and heterophily (Rogers and Bhowmik, 1970a, b). Classic GNNs, due to their fundamental mechanisms of message passing and aggregation (Battaglia et al., 2018), have demonstrated superior performance on homophilic graphs but fail to generalize to heterophily scenarios where dissimilar nodes are connected (Zhu et al., 2020; Luan et al., 2020; Liu et al., 2022). Consequently, researchers have committed to developing various structures, aiming at improving the effectiveness of GNNs on particular types of graphs (Lim et al., 2021; Liu et al., 2022; Pei et al., 2019; Luan et al., 2020). In addition to individualized designs, there are also observed recent works applying Graph Neural Architecture Search (GNAS) techniques for seeking better GNN architecture from a series of essential components (Zhou et al., 2022; Zhang et al., 2023). Despite their various routes of the existing works, these methods hold a similar methodology: Design a specialized variant of graph neural networks to tackle specific analysis tasks.
However, designing these GNN variants to work effectively in real-world applications requires extensive manual experimentation. Researchers and practitioners may engage in a meticulous process of trial and error when adjusting model architectures. This process is usually time-consuming and demands a deep understanding of the theoretical underpinnings of GNNs and the practical requirements of business. Achieving a model that works and performs efficiently and accurately in specific scenarios often involves long periods of rigorous experimentation and iterative refinement. Moreover, these GNN variants tailored for different scenarios are not significantly different from the classic GNN models theoretically (Zhu et al., 2021). In other words, though they may incorporate modifications to address specific needs or optimize performance for particular graphs, the foundational principles and underlying architectures might remain largely consistent with classic GNNs. Whether the capacities of classic GNNs can be maximized for different scenarios has become an open research problem.
In this paper, our objective is to empower classic GNNs for consistently superior performance on graphs with varying characteristics. To this end, an influential off-shelf tool, the Large Lanuange Model (LLM), is adopted to help us achieve the goal. Following the scaling law (Kaplan et al., 2020), LLMs, such as 175B-parameter GPT-3 (Brown et al., 2020), display distinct behaviors compared to smaller pre-trained language models (PLMs) like the 330M-parameter BERT (Devlin et al., 2018) and the 1.5B-parameter GPT-2 (Radford et al., 2019). LLMs exhibit surprising emergent abilities (Wei et al., 2022), enabling them to solve complex tasks that smaller models struggle with (Zhao et al., 2023a). Consequently, the research community has focused increasingly on these LLMs, acknowledging their unique and advanced capabilities (Brown et al., 2020; Touvron et al., 2023; Alayrac et al., 2022; Liu et al., 2023; Ren et al., 2023; Mysore et al., 2023).
@@
LLMs currently have not only achieved remarkable advancements in natural language processing (Brown et al., 2020; Touvron et al., 2023), but have also buoyed promising insights in other domains, such as computer vision (Alayrac et al., 2022; Liu et al., 2023, 2024; Reid et al., 2024) and recommendation systems (Ren et al., 2023; Mysore et al., 2023; Lin et al., 2024; Zhu et al., 2024), etc. This success can be attributed to the wealth of open-world knowledge stored in their large-scale parameters (Zhao et al., 2023a; Pan et al., 2024) and the emerging capabilities in logical reasoning (Huang and Chang, 2022; Brown et al., 2020). Analogously, we aim to leverage the advantages of LLMs to enhance graph machine learning, particularly the capacities of GNNs. We formulate a new paradigm that integrates LLMs with GNNs in an interactive manner, and we coin it “LLMs-as-Consultants”, following which, LLMs could explicitly contribute to the training process of GNN.
@@
Under this paradigm, we propose a framework named L O G IN , short for L LM c O nsulted G NN tra IN ing , empowering the interactive utilization of LLMs within the GNN training process. As an interactive approach, the crucial issues of L O G IN lie in what GNNs should deliver to LLMs and how to feed the LLMs’ responses back to GNNs. First, for LLMs’ inputs, we delve into prompt engineering to craft concise prompts for spotted nodes, which carry comprehensive semantic and topological information. Second, to utilize the responses from LLMs, we devise a complementary coping mechanism depending on their correctness. Specifically, compared to ground truth labels, when LLMs predict correctly, we update node features to obtain semantic enhancement. Otherwise, we impute the misclassification to the potential presence of local topological noises, hence performing structure refinement. Besides, particular criteria may serve in the selection of essential nodes to reduce time complexity. In our implementation, we adopt GNN predictive uncertainty to assess the necessity of consulting LLMs.
@@
To demonstrate the versatility and applicability of the proposed framework, we explore the effectiveness of L O G IN on node classification tasks across both homophilic and heterophilic graphs. By empirical studies, we illustrate that even classic GNNs, when employed within the proposed LLMs-as-Consultants paradigm, can achieve comparable performance to advanced GNNs with intricate designs.
@@
The remainder of this paper is organized as follows. Section 2 surveys the related research in literature. Section 3 discusses the core differences and advantages of our proposed LLMs-as-Consultants paradigm, distinguishing it from previous methods. Section 4 introduces the concepts and preliminaries of this work. Section 5 details our L O G IN approach. Section 6 illustrates the evaluation of our model and baselines in six benchmark datasets, and Section 7 concludes the paper and discusses future work.
2. Related Works
2.1. GNN Variants
Recall that early proposed GNNs (Kipf and Welling, 2016; Hamilton et al., 2017a; Velickovic et al., 2017) are primarily designed for the most typical graph-structured data, such as citation networks, which are characterized as homophilic, homogeneous, and class-balanced graphs. Here homophily means nodes with similar attributes are more likely to connect (Hamilton, 2020; Rogers and Bhowmik, 1970a). For the second one, homogeneous graphs compose different types of entities (i.e., nodes) and relations (i.e., edges) (Wang et al., 2019; Zhang et al., 2019). Class balance refers to the situation where the distribution of classes or labels within a graph is nearly even (Zhao et al., 2021; Ju et al., 2024). Consequently, when generalized to heterophilic (Bo et al., 2021; Abu-El-Haija et al., 2019; Xu et al., 2018; Chen et al., 2020), heterogeneous (Zhang et al., 2019; Wang et al., 2019; Yun et al., 2019; Fu et al., 2020), and class-imbalanced graphs (Zhao et al., 2021; Qu et al., 2021; Park et al., 2021; Liu et al., 2021; Chen et al., 2021; Song et al., 2022), classic GNNs suffer from severe performance degradation.
However, the characteristics of heterophily, heterogeneity, and class imbalance are widely recognized in the graphs constructed from real-world scenarios. For example, heterophily is evident in online transaction networks as fraudsters are more likely to form connections with normal customers (Huang et al., 2022); in molecular graphs, different types of amino acids tend to link as to compose proteins (Zhu et al., 2020). Heterogeneous graphs can represent the various entities and their relations in recommendation (Shi et al., 2018; Hu et al., 2018) and cybersecurity systems (Hu et al., 2019; Hou et al., 2017). Class imbalance is present in fraud detection based on financial transaction graphs (Liu et al., 2021; Tian et al., 2023) due to the rarity of fraudulent cases.
To address the challenges GNNs face in generalizing to a wide range of real-world applications, researchers have been motivated to tailor classic GNNs to specific scenarios, namely heterophily, heterogeneity, class imbalance, and so on.
2.1.1. GNNs Designed for Heterophily
For heterophily, since similar nodes may not tend to bond, neighbor extension and GNN architecture refinement are two major designs for variants targeting heterophilic graphs. For neighbor extension, MixHop (Abu-El-Haija et al., 2019) and H2GCN (Zhu et al., 2020) enable GNNs to aggregate information across multi-hop ego-graphs. For GNN architecture refinement, H2GCN (Zhu et al., 2020) also excludes self-loop connections for non-mixing node embedding learning to produce distinguishable node embeddings. JKNet (Xu et al., 2018) combines intermediate representations from each layer, allowing flexible adaptation of neighborhood ranges to individual nodes. H2GCN (Zhu et al., 2020) and GCNII (Chen et al., 2020) further utilize all intermediate representations and only the first layer’s node embedding at each layer with the initial residual connection to realize inter-layer combination, respectively.
2.1.2. GNNs Designed for Heterogeneity
For heterogeneity, the key to designing a specified GNN variant is to propose a suitable aggregation function (Wang et al., 2022), which can better fuse the structural and semantic information of heterogeneous graphs. To name some, HAN (Wang et al., 2019) develops a hierarchical attention mechanism to determine the significance of various nodes and meta-paths, effectively capturing both the structural and semantic information within heterogeneous graphs. HetGNN (Zhang et al., 2019) employs bi-LSTM to aggregate neighbor embeddings, facilitating the learning of deep interactions among heterogeneous nodes. GTN (Yun et al., 2019) proposes an aggregation function that can automatically discover suitable metapaths during the message-passing process. MAGNN (Fu et al., 2020) designs three main components, i.e., node content transformation for input node attributes, intra-metapath aggregation for intermediate semantic nodes, and inter-metapath aggregation to integrate messages from multiple metapaths, with the aim to enhance its expressing power of heterogeneous graphs.
2.1.3. GNNs Designed for Class Imbalance
For class imbalance, the goal of designing GNN variants is to develop a GNN classifier that effectively handles both majority and minority classes (Ju et al., 2024). For example, GraphSMOTE (Zhao et al., 2021) utilizes synthetic minority oversampling in the embedding space to boost minority class representation. ImGAGN (Qu et al., 2021) and GraphENS (Park et al., 2021) address the class imbalance by generating synthetic minority nodes. PC-GNN (Liu et al., 2021) utilizes a label-balanced sampler to pick nodes for sub-graph training and a neighborhood sampler to balance the local topology. ReNode (Chen et al., 2021) and TAM (Song et al., 2022) adjust the weights of labeled nodes in the loss function utilizing topological information.
Theoretically, the aforementioned GNN variants tailored for different scenarios are not significantly different from the classic GNNs (Zhu et al., 2021). Practically, designing these variants to work is also challenging; it requires extensive manual experimentation over a long period. Our work aims to leverage LLMs to optimize GNNs, thereby enhancing the performance of GNNs on a diverse range of graphs, ultimately matching or even surpassing these intricately designed GNN variants.
2.2. LLMs for Graphs
2.2.1. LLMs-as-Predictors
Representative methods following this paradigm, e.g. NLGraph (Wang et al., 2023a), GPT4Graph (Guo et al., 2023), GraphQA (Fatemi et al., 2023) and GraphText (Zhao et al., 2023b), attempt to harness the zero-shot and few-shot ability, along with the in-context learning ability to solve graph tasks by describing the graph structural topology in natural language. However, the complex global graph structure can hardly be compressed in a token-limited prompt, thereby utilizing LLMs solely only achieves the performance far from desired. Besides directly prompting LLMs, GraphLLM (Chai et al., 2023) conducts prefix tuning on an open-source LLM by concatenating graph-specific prefixes to the attention layers of a pre-trained LLM. InstructGLM (Ye et al., 2023) and GraphGPT (Tang et al., 2023) adapt LLMs for graph downstream tasks through instruction tuning, employing natural language and a graph-text aligner to express graph structural information, respectively. Nevertheless, neglecting the authenticated power of existing GNNs results in only moderate performance or substantial computing resource consumption.
2.2.2. LLMs-as-Enhancers
Different from the former paradigm, LLMs-as-Enhancers paradigm incorporates LLMs to enhance input graphs before GNN training. For example, TAPE (He et al., 2023) and Graph-LLM (Chen et al., 2023) prompt LLMs to interpret original texts attached to the nodes, and provide GNNs with semantically enhanced node features. SimTeG (Duan et al., 2023) parameter-efficiently finetunes (PEFT) LLMs with LoRA (Hu et al., 2021) on the textual corpus of graphs and then use the finetuned LLM to generate node representations for GNN prediction. LLM-TSE (Sun et al., 2023) explicitly instructs LLMs to produce a similarity score for two texts of two nodes, which subsequently leads to edge pruning. This methodology employs LLMs solely as a one-time data preprocessor before GNN training.
In summary, the LLMs-as-Predictors methods may result in suboptimal performance with higher resource consumption for overlooking the success of GNNs. The LLMs-as-Enhancers methods use LLMs to augment data before GNN training without integrating the LLM intelligence into GNN training. In comparison, our proposed LLM-as-Consultants paradigm aims to integrate LLMs into GNN training by consulting LLMs with a subset of nodes identified by GNNs as the most uncertain, in the attempt to enhance the classic GNNs with acceptable additional resource consumption. The core advancements of the LLMs-as-Consultants paradigm compared to previous paradigms, will be discussed detailedly in Section 3.
3. LLMs-as-Consultants Paradigm
We display the pipelines of three paradigms for integrating LLMs with graphs in Fig. 1. As shown in Fig. 1 (a) and (b), some pioneering research introduced in Section 2.2 has explored the potential of LLMs on graph data, which can be typically categorized into LLMs-as-Predictors and LLMs-as-Enhancers.
For the first kind, taking the node classification task as an example, some researchers harness the zero-shot and few-shot reasoning ability of LLMs to directly obtain node labels (Zhao et al., 2023b; Fatemi et al., 2023; Wang et al., 2023a; Guo et al., 2023). Others apply instruction tuning or prefix tuning to adapt LLMs specifically for graph tasks (Chai et al., 2023; Ye et al., 2023; Tang et al., 2023). However, most of them disregard the advantages of GNNs, and tuning LLMs may require substantial computing resources as well. For the latter one, LLMs-as-Enhancers, the LLMs are commonly utilized to enhance nodes’ semantic features (Chen et al., 2023; He et al., 2023; Duan et al., 2023) or to refine local topological structures (Sun et al., 2023). This kind of route is straightforward to apply LLMs in data preprocessing as a one-time enhancer, failing to couple the LLMs and GNNs interactively.
In contrast to existing approaches, our work introduces a novel paradigm that seamlessly integrates LLMs with GNNs through an interactive framework. We refer to this paradigm as ”LLMs-as-Consultants” (see Fig. 1 (c)). The fundamental distinction of our method lies in the recognition that LLMs can play an explicit and significant role in the training process of GNNs. This integration is not merely a superficial combination but an iteratively interactive manner where LLMs provide valuable insights and guidance that directly influence and enhance the learning capabilities of GNNs.
Figure 1. The paradigms for integrating LLMs with graphs.
@@ 4.1. Text-Attributed-Graphs (TAGs)
where $\mathcal{V}$ is a set of $N$ nodes, $\mathbf{A}\in\{0,1\}^{N\times N}$ denotes the adjacency matrix, and $s_{n}\in\mathcal{D}^{L_{n}}$ is the text attached to node $n\in\mathcal{V}$, with $\mathcal{D}$ as the word dictionary, and $L_{n}$ as the text length. Note that our proposed framework is not limited to traditional TAGs that literally have texts as original attributes. In fact, most entities and their relations can be modeled and processed as graphs, and their characteristics can be expressed in text form.
@@ 4.2. Graph Neural Networks (GNNs)
For GNN training, the texts associated with the nodes should be encoded to an embedded space. We represent the node embeddings as $\mathbf{X}\in\mathbb{R}^{N\times D}$, in which each row $\boldsymbol{}{x}_{n}$ denotes the corresponding node embedding, with $D$ as its dimension number.
A general GNN framework includes information aggregation and update processes (Gilmer et al., 2017). The aggregation function is a differentiable function for collecting and combining information from a node’s neighbors to produce a summary vector. It typically involves a form of pooling or summing the features from neighboring nodes. This process can be defined generically as follows:
(2) $\mathbf{x}_{i}^{(k)}=\text{AGGREGATE}^{(k)}\left(\left\{\mathbf{x}_{j}^{(k-1)}:j\in\mathcal{N}(i)\right\}\right)$
Note that $x_{i}^{(k)}\in\mathbb{R}^{D}$ denotes the representation of node $i$ at layer $k$, with $\mathcal{N}(i)$ as the set of neighbors of node $i$. After aggregating the features from the neighbors, the update function combines the aggregated features with the node’s own features from the previous layer. The formula for updating is as follows:
(3) $\mathbf{x}_{i}^{(k)}=\text{UPDATE}^{(k)}\left(\mathbf{x}_{i}^{(k-1)},\mathbf{x}_{i}^{(k)}\right)$
Aggregation and update are fundamental operations in GNNs, enabling nodes to exchange information and update their representations based on the graph structure and node features. These operations facilitate the learning of complex patterns and dependencies in graph-structured data. For node classification, the final node representations from the last layer are used as input to a classifier for predicting the class labels.